OpenStudy (anonymous):

Evaluate the limit \lim_{ x \rightarrow 3 } \frac { x - 3 } {x^2 + 4 x - 21 }

OpenStudy (amistre64):

well, what is the 'value' at x=3?

OpenStudy (amistre64):

if its 0/0 then we know that the linear is a factor of the quadratic, and can be removed

OpenStudy (anonymous):

its evaluate the limit..... x -> 3 x-3/x squared + 4x -21

OpenStudy (amistre64):

yeah .... the question is readable. evaluate it at x=3, if you get 0/0, then factor out the linear and youll have an equivalent function to play with

OpenStudy (anonymous):

i dont understand example if possible???

OpenStudy (amistre64):

let x=3, what are the values for top and bottom?

OpenStudy (anonymous):

\[\lim_{ x \rightarrow 3 } \frac { x - 3 } {x^2 + 4 x - 21 }\] What is the outcome when you use x=3?

OpenStudy (anonymous):

0/0

OpenStudy (amistre64):

good, then the bottom has a factor of (x-3) in it ... factor it out and cancel it

OpenStudy (anonymous):

????

OpenStudy (amistre64):

you cover polynomials before you cover rational expressions you should have already covered how to factor a quadratic

OpenStudy (anonymous):

derivative?

OpenStudy (amistre64):

no, factors (x+a)(x+b) = x^2 +(a+b)x + ab

OpenStudy (anonymous):

i know what you mean now

OpenStudy (amistre64):

:) i knew you had to have known at least at some point

OpenStudy (amistre64):

so what we have is: (x-3) -------------- (x-3)(something) which is why when x=3, the top and bottom go to zero, if we factor the bottom and cancel 1 -------------- (something) we have something that will tell us the limit since we were able to remove the offending zero

OpenStudy (anonymous):

after factoring i got 3 and -7

OpenStudy (amistre64):

those are roots, but the factored form is one step less than finding roots but fair enough x^2 + 4 x - 21 (x-3) (x+7) therefore: 1/(x+7) is our equivanelt function that we can assess the limit with when x=3, what do we get :)

OpenStudy (anonymous):

1/10

OpenStudy (anonymous):

thanks lad

OpenStudy (amistre64):

youre welcome ;)